# Turbulent gas-rich discs at high redshift: origin of thick stellar discs through 3D ‘baryon sloshing’

### v1

@@ Abstract
In response to recent observations from JWST and ALMA, we explore a new class of dynamically self-consistent models using our agama/Ramses hydrodynamic N-body simulation framework (Nexus) that mimics a plausible progenitor of the Milky Way over a wide range of disc gas fractions ($f_{\rm gas}=0,20,40,60,80,100\%$). The high gas surface densities encourage vigorous star formation, which in turn couples with the gas to drive turbulence. We show that this coupling through momentum recoil drives ‘baryon sloshing,’ i.e. a random walk of the baryonic potential minimum with respect to the centre of the total gravitational potential, $\Phi_{\rm tot}$. The amplitude of the bulk motion depends on the strength of the feedback, which in turn is directly associated with $f_{\rm gas}$. At its most extreme, when gas is the sole contributor to the disc potential ($f_{\rm gas}=100$%), the amplitude of the walk can reach up to $R\approx 5$ kpc and $|z|\approx 1$ kpc within $\Phi_{\rm tot}(R,\phi,z)$. Consistent with observations, the disc dominates over dark matter ($f_{\rm disc}\gtrsim 50$%) within $R_{s}=2.2R_{\rm disc}$, where $R_{\rm disc}$ is the exponential disc scale length. For a lower $f_{\rm disc}$ and/or $f_{\rm gas}$, the 3D sloshing amplitude and velocity are reduced. The combination of strong feedback (which unbinds the disc) and sloshing leads to the newly formed stars being dynamically heated and settling to a more spatially extended disc population compared to a settled disc with low gas fraction. The 3D heating process is roughly isotropic but its effects are more noticeable in $|z|$ due to the initial dynamical coldness of the star-forming disc. Such a disc has enhanced [$\alpha$/Fe] stellar abundances and a vertical (but no radial) gradient in stellar age and metallicity, both consistent with the Milky Way’s thick stellar disc. Contrary to earlier claims, star formation in a stationary turbulent disc does not produce thick stellar discs.
@@ 1.1 The need for new simulations
Most disc galaxies observed in the near field today went through an early phase ($z>1$) when the disc mass was dominated by gas (Tacconi et al., 2020). In order for a gas disc to support itself against collapse, high star formation rates must have been commonplace, consistent with the high gas surface densities (Genzel et al., 2010). The energetic feedback presumably coupled to the gas with moderate efficiency ($\sim 10$%) and provided a natural source of (turbulent) pressure to sustain the gas-rich medium (Agertz et al., 2021; Bland-Hawthorn et al., 2024). This was an intensely energetic era that can be traced to $z\sim 10$ by our most powerful telescopes, including the Atacama Large Millimeter Array (ALMA) and the James Webb Space Telescope (JWST). Today, Milky Way analogues have a low gas fraction ($f_{\rm gas}\lesssim 10$%), and so provide limited information on the gas processes that were active at those early times. Here, we seek to gain some physical insight about the nature of turbulent gas-rich discs at high redshift (Fig. 1) through controlled simulations made possible with the Nexus framework (Tepper-Garcia et al., 2024).
Figure 1: Edge-on gas surface density maps taken from our early disc simulations at $t_{\rm o}=1\,{\rm Gyr}$. We adopt $f_{\rm disc}=0.5$ for the Milky Way progenitor at $z\sim 3$. The top panels are for five different initial gas fractions (as indicated) with no halo accretion; the bottom five are the matching simulations that include halo accretion from cooling hot gas. From left to right, for both rows, the galaxy’s centre falls at the RHS edge in the first four panels, but is centred in the last panel, where the spatial scale is indicated. The extent of the disc-halo interaction is a reflection of the disc-wide star formation shown in Fig. 2.
Our new work is in response to mounting evidence for gas-dominant, star-forming discs in the redshift range $z\sim 1-8$ (Chapman et al., 2004; Genzel et al., 2006; Förster Schreiber et al., 2006; Shapiro et al., 2008a; Swinbank et al., 2012; Hodge et al., 2019; Tsukui & Iguchi, 2021; Rizzo et al., 2022; Rowland et al., 2024; Roman-Oliveira et al., 2024; Tsukui et al., 2024b). There are clear cases where the gas entirely dominates the baryon disc fraction (e.g. Roman-Oliveira et al., 2024; Rowland et al., 2024; Amvrosiadis et al., 2024; Hodge et al., 2024). To date, there has been limited work on realistic high-resolution simulations of isolated gas-rich galaxies sustained by underlying star formation and turbulence (e.g. van Donkelaar et al., 2022a). Here we provide a sophisticated simulation framework (Nexus; see Tepper-Garcia et al. 2024) to explore gas-dominant disc systems within the context of multi-component galaxy models (stellar disc and bulge, live dark matter halo, hot gaseous corona, multiphase ISM with star formation and metal production, etc.). Thus, it is now possible to explore disc evolution over the full range of $f_{\rm gas}=0-100\%$ in controlled experiments, as illustrated in Fig. 1.
@@ 1.2 A new era in disc research
The formation of galactic discs has been one of the fastest developing areas of galactic astronomy in recent years. A host of observational results is revolutionising this field of research, including the discovery of dynamically cold galactic discs at very high redshift ($81<z<$ 8) with ALMA and JWST (e.g. Ferreira et al., 2022; Nelson et al., 2023; Ferreira et al., 2023; Kartaltepe et al., 2023; Pope et al., 2023), the existence of already-mature discs containing bars (Costantin et al., 2023a; Guo et al., 2024) and spiral structures (Kuhn et al., 2024) as early as 2 Gyr after the Big Bang. These discoveries concern objects that are expected to be massive, settling to discs earlier than Milky Way-like galaxies (disc downsizing, see Tsukui et al. 2024a), in some cases evolving into ellipticals in the present-day Universe. Exactly when Milky Way-like galaxies began to build their discs is uncertain (Mo et al., 1998; Santistevan et al., 2020). Pope et al. (2023) present a galaxy at redshift 4.3 dominated by a stellar disc ($V_{\rm circ}/\sigma_{\rm LOS}\approx 5$) with baryon mass $M_{\star}\approx 2\times 10^{9}$ M, a value in line with the estimated stellar mass of Milky Way (MW) precursors at such epochs (van Dokkum et al., 2013).
In our own Galaxy, the oldest and most metal-poor stars with disc kinematics are part of the thick disc and its metal-poor extension. The thick disc is traditionally defined in terms of its warm kinematics and its enhanced [$\alpha$/Fe] abundances compared to the solar value; this population is largely confined to $R<10$ kpc (Hayden et al., 2015). Thick disc stars are thus expected to hold the key to understanding and constraining the physical mechanisms that govern how and when these dissipative structures formed. Unfortunately, the epoch and duration of formation of this population cannot be determined precisely from stellar ages, even with recent advances in asteroseismology (Sharma et al., 2019). Some studies suggest thick disc formation lasted for 2–3 Gyr (Haywood et al., 2013a; Snaith et al., 2015), while others suggest a significantly faster formation, on the order of 1 Gyr (Miglio et al., 2021). In all cases, the bulk of the stellar mass of this population was in place 10 Gyr ago (at $z\approx 2$), with $M_{\star}$ in the range 20% (Bland-Hawthorn & Gerhard, 2016) to 50% (Snaith et al., 2014) of the total present-day stellar mass of the disc.
@@
After the first structures were identified in high-redshift discs (e.g. Elmegreen & Elmegreen, 2005; Genzel et al., 2008; Shapiro et al., 2008b), the idea that massive gas clumps could provide a means of thickening the stellar component by diffusion of orbits was proposed by Bournaud et al. (2009). As simulations of primordial discs improved, the proposal that thick discs could arise from stars forming in a thick turbulent ISM could finally be tested. Other work suggested that stars forming in a turbulent, kinematically hot ISM generate a thick stellar disc, which can be further heated by mergers or interactions with satellites (Stinson et al., 2013; Ma et al., 2017; Bird et al., 2021; van Donkelaar et al., 2022b). Here we challenge this claim by showing that the effect is not seen in higher-resolution Nexus simulations. Another scenario comes from Meng & Gnedin (2021) who, using cosmological simulations, find that it is the accumulation of the successive generations of stars forming in thin discs of different orientations that gradually produces a thick disc.
@@
Previous scenarios have generally focused on providing a way to thicken a disc, or have assumed the stars are born thick. However, the MW thick disc has important properties that were recently discovered, particularly from large spectroscopic surveys. First, the scale height of the thick disc is not constant, but increases from the most metal-rich and least $\alpha$-enhanced stars to the most metal-poor and most $\alpha$-enhanced (the so-called mono-abundance populations, see Bovy et al. 2012b; Mackereth et al. 2017a). The $\alpha$-rich disc is radially compact, with a scale length of about 2 kpc which seems to be independent of the mono-abundant population (see Bovy et al., 2012b, their Fig. 5). It is unclear whether the thick disc has a synchronized, monotonic, chemical evolution, with little radial dependence (Haywood et al., 2013a, 2015) or, conversely, an inside-out formation leading to a (steep) radial metallicity gradient (e.g. Ratcliffe et al., 2024). The former case is expected to lead to a tight age-metallicity relation, as observed in several studies (Haywood et al., 2013a; Xiang & Rix, 2022a; Casamiquela et al., 2024), while the latter implies uncorrelated age-chemistry distributions (Miglio et al., 2021; Ciucă et al., 2021; Anders et al., 2023). All these features constitute further constraints for any model of early disc formation in the Milky Way. The aim of our paper is to revisit some of these ideas and to explore a radically new mechanism for disc thickening.
@@
The structure of the paper is as follows. We summarize the Nexus models in Section 2 and provide links to the simulation movies. In Section 3, we discuss the physical processes seen in the simulations, with a focus on the processes that lead to baryon sloshing. In Section 4, we ask whether the rapid growth of instabilities seen in the simulations are supported by theory. In Section 5, we discuss the implications of our new work and suggest future studies based on kinematic and morphological signatures. In Sec. 6, we place our new work in context and look forward to future improvements in the Nexus framework.
2 Galaxy simulations in brief
@@ 2.1 Models and movies
In earlier papers, we presented our new high-resolution (parsec scale) simulation framework (Nexus), which is ideally suited to studying how turbulent discs evolve dynamically (Tepper-Garcia et al., 2024). These ecosystems manifest intermittent and long-lived behaviour, including disc-halo interaction, bars and spiral arms, and bulge formation (Bland-Hawthorn et al., 2024). We were able to generate “bar-like” phenomena for the first time at all gas fractions, even in fully gas-dominated turbulent discs, including a remarkable “radial shear flow” with a characteristic signature that sets up before the central bulge is formed. As for the stellar discs, these bar-like signatures only arise in discs that dominate the local gravitational potential (Bland-Hawthorn et al., 2023).
The choice of parameters for our progenitor analogues is described in Bland-Hawthorn et al. (2024). This present work is focussed on objects with halo masses of order 1011 M, a dynamical mass that is characteristic of the recent JWST discoveries at $z\sim 1-6$ (Guo et al., 2022; Le Conte et al., 2023; Costantin et al., 2023b). This is the expected mass of a Milky Way progenitor at $z\approx 3$ (Fig. 1 in Bland-Hawthorn & Gerhard, 2016). In high-redshift galaxies, the observed high gas fraction and enhanced gas surface densities (Tacconi et al., 2020) lead to elevated star-formation rates, and the output energy and momentum couple with the gas to drive the turbulence (Agertz et al., 2009; Jiménez et al., 2023). The adopted star/galaxy formation physics is presented in Agertz et al. (2013) and Agertz et al. (2021). The star formation activity feeds back sufficient energy and momentum to maintain the turbulent support. (We note, in passing, that turbulent gas-rich systems have also been explored where the energy injection arises from self gravity rather than star formation and feedback; see Fensch et al. 2023.)
$M_{\star}$ $M_{\rm gas}$ $R_{\rm disc}$ $f_{\rm disc}$ $f_{\rm gas}$ Label High Low
@@
Table 1: Overview of galaxy models. The DM host halo properties ($M_{\rm halo}=10^{11}$ M; $R_{\rm vir}=37$ ckpc; $r_{\rm s}=9.2$ ckpc) were identical across models, and they are consistent with a MW-progenitor analogue at $z\approx 3$. Table columns are as follows: (1) pre-existing disc stellar mass; (2) disc gas mass; (3) disc scale length; (4) disc-to-total mass ratio (Eq. 1); (5) Gas to total disc mass fraction at $t=0$ (Eq. 2) -- the double brackets indicate accreting gas from a hot corona with a mass $M\approx 5.5\times 10^{9}$ M at $t=0$; note that the hot halo mass is not accounted for in $f_{gas}$. (6) Model designation; (7,8) No. of high and low resolution models with different random seeds.
Our first step is to consider an isolated galactic ecosystem in dynamical equilibrium, both with and without smooth accretion from the ambient hot corona. All computations were carried out with the Ramses N-body/hydrodynamics code (Teyssier, 2002) at low ($N_{\rm lo}\sim 10^{7}$ elements) and high ($N_{\rm hi}\sim 10^{8}$ elements) resolution, including star formation and metal production (see Tab. 1). In defining a Milky Way progenitor, we adopted a model with three key components: a live dark matter halo, a massive stellar/gaseous disc, and a hot coronal gas filling the live dark matter halo, which serves to supply the disc with a smooth flow of accreting gas after it cools.
@@
Here, $V_{c}(R)$ is the circular velocity at a radius $R$, $R_{\rm disc}$ is the exponential disc scale length, and $R_{e}=2.2R_{\rm disc}$ is the traditional scale length adopted in studies of discs. (ii) The gas mass fraction within the same radial scale:
$f_{\rm gas}=\left(\frac{M_{\rm disc,gas}}{M_{\rm disc}}\right)_{R_{e}=2.2R_{\rm disc}},\,$ (2)
@@
The values for all parameters are given in Table 1. The simulations are grouped into different categories -- low resolution, high resolution, and repeated cases with different random seeds. There were 6 types of simulations distinguished by the different initial gas fractions ($f_{\rm gas}=0,20,40,60,80,100\%$); these are run at low resolution for a total of 2 Gyr because we are focussed on the high-redshift universe ($z\gtrsim 3$). All movies are available at our website http://www.physics.usyd.edu.au/turbo_discs/. The filename convention is fdXX_fgYY_nac for “no halo accretion” models, and fdXX_fgYY_ac for “halo accretion” models. Here, XX is the disc mass fraction percentage and YY is the gas fraction percentage. If the simulation had multiple versions generated by different random seeds, the file is referred to as fdXX_fgYY_nac.ZZ, where ZZ is the version number.
In all models, the gas fraction declines as more of the mass is locked up in stars (see Fig. 3 in Bland-Hawthorn et al., 2024). For models with $f_{\rm gas}<100\%$, we included a pre-existing (‘old’) stellar disc such that $f_{\rm old}=1-f_{\rm gas}$. For models with $f_{\rm gas}\approx 100\%$, we include a tiny pre-existing stellar population ($f_{\rm old}\approx 1\%$) as a useful tracer of the galaxy’s gas-driven evolution. The baryon mass was preserved in both the accreting and non-accreting halo models. In the former case, the total baryon mass of the disc increases with time. Note that the initial scale length of the disc (both gas and stars) was roughly maintained at $R_{\rm disc}\approx 1.8$ kpc across models. The above parameters are broadly consistent with arguably the best Milky Way progenitor analogue to date, the object CEERS-2112 at a photometric redshift of $z\approx 3.0$ (Costantin et al., 2023b).

### v2

@@ Abstract
In response to recent observations from JWST and ALMA, we explore a new class of dynamically self-consistent models that mimics a plausible progenitor of the Milky Way over a wide range of disc gas fractions, $f_{\rm gas}$. The high gas surface densities encourage vigorous star formation, which in turn couples with the gas to drive turbulence. We show that this coupling through momentum recoil drives a random walk of the baryonic potential minimum with respect to total gravitational potential, $\Phi_{\rm tot}(R,\phi,z)$. The amplitude of the bulk motion depends on the feedback strength, which in turn is directly associated with $f_{\rm gas}$. At its most extreme, when gas is the sole contributor to the disc potential ($f_{\rm gas}=100$%), the amplitude of the walk can reach up to $R\approx 5$ kpc within $\Phi_{\rm tot}$. The disc dominates over dark matter ($f_{\rm disc}\gtrsim 50$%) within $R_{s}=2.2R_{\rm disc}$, where $R_{\rm disc}$ is the exponential disc scale length. For a lower $f_{\rm disc}$ and/or $f_{\rm gas}$, the 3D sloshing amplitude and velocity are reduced. The combination of strong feedback and sloshing leads to the newly formed stars being dynamically heated and settling to a more spatially extended disc population. The 3D heating process is roughly isotropic but its effects are more noticeable in $|z|$ due to the initial dynamical coldness of the star-forming disc. Such a disc has enhanced $\alpha$/Fe] stellar abundances and a vertical (but no radial) gradient in stellar age and metallicity, both consistent with the Milky Way’s thick stellar disc.
@@ 1.1 The need for new simulations
Most disc galaxies observed in the near field today went through an early phase ($z>1$) when the disc mass was dominated by gas (Tacconi et al., 2020). In order for a gas disc to support itself against collapse, high star formation rates must have been commonplace, consistent with the high gas surface densities (Genzel et al., 2010). The energetic feedback presumably coupled to the gas with moderate efficiency ($\sim 10$%) and provided a natural source of (turbulent) pressure to sustain the gas-rich medium (Agertz et al., 2021; Bland-Hawthorn et al., 2024). This was an intensely energetic era that can be traced to $z\sim 10$ by our most powerful telescopes, including the Atacama Large Millimeter Array (ALMA) and the James Webb Space Telescope (JWST). Today, Milky Way analogues have a low gas fraction ($f_{\rm gas}\lesssim 10$%), and so provide limited information on the gas processes that were active at those early times. Here, we seek to gain some physical insight about the nature of turbulent gas-rich discs at high redshift through controlled simulations made possible with the Nexus framework (Tepper-García et al., 2024).
Our new work is in response to mounting evidence for gas-dominant, star-forming discs in the redshift range $z\sim 1-8$ (Chapman et al., 2004; Genzel et al., 2006; Förster Schreiber et al., 2006; Shapiro et al., 2008; Swinbank et al., 2012; Hodge et al., 2019; Tsukui & Iguchi, 2021; Rizzo et al., 2022; Rowland et al., 2024; Roman-Oliveira et al., 2024; Tsukui et al., 2024b). There are clear cases where the gas entirely dominates the baryon disc fraction (e.g. Roman-Oliveira et al., 2024; Rowland et al., 2024; Amvrosiadis et al., 2025; Hodge et al., 2025). To date, there has been limited work on realistic high-resolution simulations of isolated gas-rich galaxies sustained by underlying star formation and turbulence (e.g. van Donkelaar et al., 2022). Here we provide a sophisticated simulation framework (Nexus; see Tepper-García et al. 2024) to explore gas-dominant disc systems within the context of multi-component galaxy models (stellar disc and bulge, live dark matter halo, hot gaseous corona, multiphase ISM with star formation and metal production, etc.). Thus, it is now possible to explore disc evolution over the full range of $f_{\rm gas}=0-100\%$ in controlled experiments, as illustrated in Fig. 1. (The parameters $f_{\rm disc}$ and $f_{\rm gas}$ are defined in Sec. 2.)
@@ 1.2 A new era in disc research
The formation of galactic discs has been one of the fastest developing areas of galactic astronomy in recent years. A host of observational results is revolutionising this field of research, including the discovery of dynamically cold galactic discs at very high redshift ($1<z<8$) with ALMA and JWST (e.g. Ferreira et al., 2022; Nelson et al., 2024; Ferreira et al., 2023; Kartaltepe et al., 2023; Pope et al., 2023), the existence of already-mature discs containing bars (Costantin et al., 2023; Ratcliffe et al., 2025) and spiral structures (Kuhn et al., 2024) as early as 2 Gyr after the Big Bang. These discoveries concern objects that are expected to be massive, settling to discs earlier than Milky Way-like galaxies (disc downsizing, see Tsukui et al. 2024a), in some cases evolving into ellipticals in the present-day Universe. Exactly when Milky Way-like galaxies began to build their discs is uncertain (Mo et al., 1998; Santistevan et al., 2020). Pope et al. (2023) present a galaxy at redshift 4.3 dominated by a stellar disc ($V_{\rm circ}/\sigma_{\rm LOS}\approx 5$) with baryon mass $M_{\star}\approx 2\times 10^{9}$ M, a value in line with the estimated stellar mass of Milky Way (MW) precursors at such epochs (van Dokkum et al., 2013).
In our own Galaxy, the oldest and most metal-poor stars with disc kinematics are part of the thick disc and its metal-poor extension. The thick disc is traditionally defined in terms of its warm kinematics and its enhanced $\alpha$/Fe] abundances compared to the solar value; this population is largely confined to $R<10$ kpc (Hayden et al., 2015). Thick disc stars are thus expected to hold the key to understanding and constraining the physical mechanisms that govern how and when these dissipative structures formed. Unfortunately, the epoch and duration of formation of this population cannot be determined precisely from stellar ages, even with recent advances in asteroseismology (Sharma et al., 2019). Some studies suggest thick disc formation lasted for 2–3 Gyr (Haywood et al., 2013; Snaith et al., 2015), while others suggest a significantly faster formation, on the order of 1 Gyr (Miglio et al., 2021). In all cases, the bulk of the stellar mass of this population was in place 10 Gyr ago (at $z\approx 2$), with $M_{\star}$ in the range 20% (Bland-Hawthorn & Gerhard, 2016) to 50% (Snaith et al., 2014) of the total present-day stellar mass of the disc.
@@
After the first structures were identified in high-redshift discs (e.g. Elmegreen & Elmegreen, 2005; Genzel et al., 2008; Shapiro et al., 2008), the idea that massive gas clumps could provide a means of thickening the stellar component by diffusion of orbits was proposed by Bournaud et al. (2009). These authors carried out sticky particle simulations rather than more realistic hydrodynamic simulations. As simulations of primordial discs improved (Dekel et al., 2009), the proposal that thick discs could arise from stars forming in a thick turbulent ISM could finally be tested. Other work suggested that stars forming in a turbulent, kinematically hot ISM generate a thick stellar disc, which can be further heated by mergers or interactions with satellites (Stinson et al., 2013; Ma et al., 2017; Bird et al., 2021; van Donkelaar et al., 2022). Here we challenge this claim by showing that the effect is not seen in higher-resolution Nexus simulations. Another scenario comes from Meng & Gnedin (2021) who, using cosmological simulations, find that it is the accumulation of the successive generations of stars forming in thin discs of different orientations that gradually produces a thick disc.
@@
Previous scenarios have generally focused on providing a way to thicken a disc, or have assumed the stars are born thick. However, the MW thick disc has important properties that were recently discovered, particularly from large spectroscopic surveys. First, the scale height of the thick disc is not constant, but increases from the most metal-rich and least $\alpha$-enhanced stars to the most metal-poor and most $\alpha$-enhanced (the so-called mono-abundance populations, see Bovy et al. 2012b; Mackereth et al. 2017). The $\alpha$-rich disc is radially compact, with a scale length of about 2 kpc which seems to be independent of the mono-abundant population (see Bovy et al., 2012b, their Fig. 5). It is unclear whether the thick disc has a synchronized, monotonic, chemical evolution, with little radial dependence (Haywood et al., 2013, 2015) or, conversely, an inside-out formation leading to a (steep) radial metallicity gradient (e.g. Ratcliffe et al., 2025). The former case is expected to lead to a tight age-metallicity relation, as observed in several studies (Haywood et al., 2013; Xiang & Rix, 2022; Casamiquela et al., 2024), while the latter implies uncorrelated age-chemistry distributions (Miglio et al., 2021; Anders et al., 2023). All these features constitute further constraints for any model of early disc formation in the Milky Way. The aim of our paper is to revisit some of these ideas and to explore a new mechanism for disc thickening.
@@
The structure of the paper is as follows. In Section 2, we show how controlled simulations link to cosmological simulations, summarize the Nexus models and provide links to the simulation movies. In Section 3, we discuss the physical processes seen in the simulations, with a focus on the processes that lead to baryon sloshing. In Section 4, we ask whether the rapid growth of instabilities seen in the simulations are supported by theory. In Section 5, we discuss the implications of our new work and suggest future studies based on kinematic and morphological signatures. In Sec. 6, we place our new work in context and look forward to future improvements in the Nexus framework.
Figure 1: Edge-on gas surface density maps taken from our early disc simulations at $t_{\rm o}=1\,{\rm Gyr}$. We adopt $f_{\rm disc}=0.5$ for the Milky Way progenitor at $z\sim 3$. The top panels are for five different initial gas fractions (as indicated) with no halo accretion; the bottom five are the matching simulations that include halo accretion from cooling hot gas. From left to right, for both rows, the galaxy’s centre falls at the right-hand side (RHS) edge in the first four panels, but is centred in the last panel, where the spatial scale is indicated. The extent of the disc-halo interaction is a reflection of the disc-wide star formation shown in Fig. 2.
2 Nexus simulations in brief
2.1 Controlled simulations in the context of cosmological N-body simulations
In a seminal paper, Mo et al. (1998) introduced a simple theoretical framework describing how galaxy discs form from cooling hot gas within virialized dark matter halos, with their properties largely governed by angular momentum conservation (see also the updated treatment in Stern et al. (2020)). This work laid the foundation for semi-analytic models (SAMs) of galaxy formation. An unrecognized assumption in these models is that the circumgalactic medium (CGM) becomes structurally and dynamically stable (i.e., virialized) out to the virial radius at some early epoch. However, this assumption cannot hold in a strict sense, as the models also require that hot gas continues to cool within some inner radius.
Modern cosmological simulations diverge significantly in their predictions of CGM properties for Milky Way-like galaxies, mostly because of different implementations of star-formation driven galactic winds (q.v. Stern et al., 2021; Wright et al., 2024). Controlled simulations are superior to cosmological simulations in tracing the evolution of galactic winds; the former include more realistic processes at much higher spatial resolution (q.v. Sike et al., 2025). Observationally, absorption-line studies of H I and heavy elements along quasar sightlines suggest that the CGM began to settle into a more stable state by around $z\gtrsim 2$ (Chen, 2012), consistent with CGM models showing self-similar behavior since $z\sim 4$ (Pallottini et al., 2014). This agreement supports our working assumption for comparisons with controlled simulations, such as those in the VINTERGATAN suite (Agertz et al., 2021). Specifically, we adopt characteristic CGM properties from VINTERGATAN progenitors at $z\approx 3$, a redshift by which the CGM appears largely stabilized (Rufo-Pastor et al., in prep.). In future work, we will compare the evolutionary trajectories produced by these complementary approaches to modelling galaxy evolution.
@@ 2.2 Models and movies
In earlier papers, we presented our new high-resolution (parsec scale) simulation framework (Nexus), which is ideally suited to studying how turbulent discs evolve dynamically (Tepper-García et al., 2024). These ecosystems manifest intermittent and long-lived behaviour, including disc-halo interaction, bars and spiral arms, and bulge formation (Bland-Hawthorn et al., 2024). We were able to generate “bar-like” phenomena for the first time at all gas fractions, even in fully gas-dominated turbulent discs, including a remarkable “radial shear flow” with a characteristic signature that sets up before the central bulge is formed. As for the stellar discs, these bar-like signatures only arise in discs that dominate the local gravitational potential (Bland-Hawthorn et al., 2023).
The choice of parameters for our progenitor analogues is described in Bland-Hawthorn et al. (2024). This present work is focussed on objects with halo masses of order 1011 M, a dynamical mass that is characteristic of the recent JWST discoveries at $z\sim 1-6$ (Guo et al., 2023; Le Conte et al., 2024; Costantin et al., 2023). This is the expected mass of a Milky Way progenitor at $z\approx 3$ (Fig. 1 in Bland-Hawthorn & Gerhard, 2016). In high-redshift galaxies, the observed high gas fraction and enhanced gas surface densities (Tacconi et al., 2020) lead to elevated star-formation rates, and the output energy and momentum couple with the gas to drive the turbulence (Agertz et al., 2009; Jiménez et al., 2023). The adopted star/galaxy formation physics is presented in Agertz et al. (2013) and Agertz et al. (2021). The star formation activity feeds back sufficient energy and momentum to maintain the turbulent support. (We note, in passing, that turbulent gas-rich systems have also been explored where the energy injection arises from self gravity rather than star formation and feedback; see Fensch et al. 2023.)
$M_{\star}$ $M_{\rm gas}$ $R_{\rm disc}$ $f_{\rm disc}$ $f_{\rm gas}$ Label High Benchmark
@@
Table 1: Overview of galaxy models of the initial parameters. The DM host halo properties ($M_{\rm halo}=10^{11}$ M; $R_{\rm vir}=37$ ckpc; $r_{\rm s}=9.2$ ckpc) are identical across models, and they are consistent with a MW-progenitor analogue at $z\approx 3$. Table columns are as follows: (1) pre-existing disc stellar mass (vertical scale height == 250 pc for all models); (2) disc gas mass; (3) disc scale length; (4) disc-to-total mass ratio (Eq. 1); (5) Gas to total disc mass fraction at $t=0$ (Eq. 2) - the double brackets indicate accreting gas from a hot corona with a mass $M\approx 5.5\times 10^{9}$ M at $t=0$; note that the hot halo mass is not accounted for in $f_{gas}$. (6) Model designation; (7,8) No. of high and benchmark resolution models with different random seeds.
Our first step is to consider an isolated galactic ecosystem in dynamical equilibrium, both with and without smooth accretion from the ambient hot corona. All computations were carried out with the Ramses N-body/hydrodynamics code (Teyssier, 2002) at benchmark ($N_{\rm lo}\sim 10^{7}$ elements) and high ($N_{\rm hi}\sim 10^{8}$ elements) resolution, including star formation and metal production (see Tab. 1); the number of effective gas “cells” is a factor of 10 higher in both cases. In defining a Milky Way progenitor, we adopted a model with three key components: a live dark matter halo, a massive stellar/gaseous disc, and a hot coronal gas filling the live dark matter halo, which serves to supply the disc with a smooth flow of accreting gas after it cools.
@@
Here, $V_{c}(R)$ is the circular velocity at a radius $R$, $R_{\rm disc}$ is the exponential disc scale length, and $R_{e}=2.2R_{\rm disc}$ is the traditional scale length adopted in studies of discs as it corresponds approximately to the “turnover radius” in the rotation curve [^1] This radius contains about two-thirds of the disc’s total mass.; (ii) The gas mass fraction within the same radial scale:
$f_{\rm gas}=\left(\frac{M_{\rm disc,gas}}{M_{\rm disc}}\right)_{R_{e}=2.2R_{\rm disc}}$ (2)
@@
The values for all parameters are given in Table 1. The simulations are grouped into different categories - benchmark resolution, high resolution, and repeated cases with different random seeds. There were 6 types of simulations distinguished by the different initial gas fractions ($f_{\rm gas}=0,20,40,60,80,100\%$); these are run at benchmark resolution for a total of 2 Gyr because we are focussed on the high-redshift universe ($z\gtrsim 3$).
In all models, the gas fraction declines as more of the mass is locked up in stars (see Fig. 3 in Bland-Hawthorn et al., 2024). For models with $f_{\rm gas}<100\%$, we included a pre-existing (‘old’) stellar disc such that $f_{\rm old}=1-f_{\rm gas}$. For models with $f_{\rm gas}\approx 100\%$, we include a tiny pre-existing stellar population ($f_{\rm old}\approx 1\%$) as a useful tracer of the galaxy’s gas-driven evolution. The total mass of stars produced in 2 Gyr from low to high gas fraction spans $2.6-6.2\times 10^{9}$ M for the accreting halo models, and $0.6-4.0\times 10^{9}$ M for the non-accreting models. The baryon mass was preserved in both the accreting and non-accreting halo models. In the former case, the total baryon mass of the disc increases with time. Note that the initial scale length of the disc (both gas and stars) was roughly maintained at $R_{\rm disc}\approx 1.8$ kpc across models. The above parameters are broadly consistent with arguably the best Milky Way progenitor analogue to date, the object CEERS-2112 at a photometric redshift of $z\approx 3.0$ (Costantin et al., 2023).
In Table 1, the filename convention is fdXX_fgYY_nac for “no halo accretion” models, and fdXX_fgYY_ac for “halo accretion” models. Here, XX is the disc mass fraction percentage and YY is the gas fraction percentage. If the simulation has multiple versions generated by different random seeds, the file is referred to as fdXX_fgYY_nac.ZZ, where ZZ is the version number. At our website, [^2] http://www.physics.usyd.edu.au/turbo_discs/ we provide a series of movies (animations) that show the evolution of each of our model galaxies. There are two types of animations: 1) ‘on-the-fly’ animations; and 2) ‘post-processed’ animations. The ‘on-the-fly’ animations were created at simulation runtime; these have a high time resolution ($\delta t\approx 1$ Myr), and they show the evolution of the newly formed stars, of the gas density, and of the gas temperature on a face-on projection. The ‘post-processed’ animations have been created from the simulation outputs; they have a lower time resolution ($\delta t\approx 10$ Myr), and they show the evolution of the surface density of all disk components (gas, pre-existing stars, and - if available - newly formed stars, gas) along three orthogonal projections.